The internal pH of a muscle is 6.8. Calculate the [H2PO4-]/[HPO42-] ratio in the cell. The second dissociation constant of phosphoric acid is 6.31x10-8

1
Expert's answer
2022-02-25T15:17:04-0500

Solution:

H2PO4 ⇌ H+ + HPO42−, K2 = 6.31×10−8

pK2 = −log(K2) = −log(6.31×10−8) = 7.1999 = 7.20


Henderson–Hasselbalch equation can be used.

pH = pKa + log([A] / [HA])


pH = 6.80

pKa = pK2 = 7.20

[A] = HPO42−

[HA] = H2PO4

Therefore,

6.80 = 7.20 + log([HPO42−] / [H2PO4])

−0.40 = log([HPO42−] / [H2PO4])

[HPO42−] / [H2PO4] = 10−0.40 = 0.398

[H2PO4] / [HPO42−] = 1 / 0.398 = 2.51

[H2PO4] / [HPO42−] = 2.51


Answer: The ratio [H2PO4] / [HPO42−] is 2.51